michaela2006:

Ernie earned $121.50 more than Kaylin. Tria earned $35 less than 3 times the amount Kaylin earned. Ernie, Tria, and Kaylin earned $579 in all. How much did Tria earn?

jhonyy9:

Ernie earned $121.50 more than Kaylin. Tria earned $35 less than 3 times the amount Kaylin earned. Ernie, Tria, and Kaylin earned $579 in all. How much did Tria earn? so note what Ernie earned with E what Kaylin earned with K what Tria earned with T so Ernie earned $121,50 +K so E = 121 +K Tria earned 3*K -35 so T = 3*K -35 E + T + K = 579 using these above wrote can you calcule the correct answers ?

jhonyy9:

ohh @michaela2006 is offline so just solve this like a system of equations with 3 variables

jhonyy9:

E = 121 +K T = 3K -35 E + T + K = 379

jhonyy9:

so how you solve this ? any idea ?

jhonyy9:

so the easyst way to solve is to substitute the first two equations in the 3rd and calcule the value of K yes ?